# A Simple Knowledge Distillation Framework for Open-world Object Detection

### v1

@@ Introduction
Down-Weight Training Strategy
where the uses the common regression loss which consists of box and GIOU loss (Rezatofighi et al. 2019). and represent the box score and classification loss, respectively. They all leverage the common sigmoid focal loss (Lin et al. 2017). For simplicity, we omit them. In addition, the , , and all utilize correspondingly down-weight loss function we propose, the formulations are shown as follows:
Inference
Experiment
Datasets and Metrics
For a fair comparison, we implement the experiments on two mainstream splits of MS-COCO (Lin et al. 2014), and Pascal VOC (Everingham et al. 2010) dataset. We group the classes into a set of non-overlapping tasks . The class in task only appears in tasks where . In task , classes encountered in and are considered as known and unknown classes, respectively.
OWOD SPLIT (Joseph et al. 2021) splits the 80 classes of MS-COCO into 4 tasks and selects a training set for each task from the MS-COCO and Pascal VOC training set. Pascal VOC testing and MS-COCO validation set are used for evaluation.
MS-COCO SPLIT (Gupta et al. 2022) mitigates data leakage across tasks in (Joseph et al. 2021) and is more challenging. The training and testing data are selected from MS-COCO.
Metrics: Following the most commonly used evaluation metric for object detection, we use mean average precision (mAP) to evaluate the known objects. Following OW-DETR (Gupta et al. 2022), U-Recall which measures the ability of the model to retrieve unknown object instances for OWOD problems, is used as the metric for unknown objects.
Implementation Details
The multi-scale feature extractor consists of a Resnet-50(He et al. 2016) pretrained on ImageNet(Deng et al. 2009) in a self-supervised(Caron et al. 2021) manner and a deformable transformer encoder whose number of layers is set to 6. For the two cascade decoders, we all use the deformable transformer decoder, and the number of layers is set to 6, too. We set the number of queries , the dimension of the embeddings , and the number of pseudo-labels . We set GLIP (Li et al. 2022) as the large pre-trained vision-language grounding model and categories of LVIS dataset (Gupta, Dollar, and Girshick 2019) as text prompts to assist the training process. For GLIP, we use the GLIP-L (Li et al. 2022) without finetuning on the COCO dataset (Lin et al. 2014). It consists of Swin-Large (Liu et al. 2021), text encoder of CLIP (Radford et al. 2021), DyHead (Dai et al. 2021), BERT Layer (Devlin et al. 2018) and Fusion Module (Li et al. 2022). For the incremental object detection experiments, we only use our open-world detector without the help of GLIP.
Ablation Study
In this subsection, a set of experiments are designed to clearly understand the contribution of each of the constituent components. We conducted all experiments on the OWOD split.
| Task IDs | Task 1 | Task 2 | Task 3 | |||
| Inference | Unknown | |||||
| Param# | Rate | Recall() | ||||
| SPLIT | Metrics | (s/img) | ||||
| OWOD | GLIP | 321.9M | 9.22 | 37.0 | 35.5 | 34.9 |
| Ours | 42.9M | 0.08 | 39.0 | 36.7 | 36.1 | |
| MS-COCO | GLIP | 321.9M | 9.22 | 52.6 | 54.5 | 53.3 |
| Ours | 42.9M | 0.08 | 60.9 | 60.0 | 58.6 | |
@@
Ablating components: To study the contribution of each component, we design ablation experiments in Table.1 (a). We set GLIP as our baseline, the Distillation improves the performance on unknown object detection but reduce the detector’s ability for known objects. Adding the down-weight training loss function significantly improves the performance on detecting known objects and incremental object detecting, achieving significant absolute gains up to more than 10 points. As we have analyzed, the cascade decoupling structure alleviates the inclusion of unknown objects on the known detecting performance and the confusion between the categories and locations of the same objects. It significantly improves the performance of detecting known objects with absolute gains up to more than 10 points, too. What’s more, these two combine effectively, improving performance with absolute gains up to 15.7, 21.1 and 16.0, without reducing the ability to detect unknown objects. Thus, each component has a critical role to play in open-World object detection.
@@
Ablating different prompts: To ablate the impact of text prompts for knowledge distillation, we do experiments on different text prompts, i.e.LVIS, LVIS without COCO and COCO as shown in Table.1(a). From the experiment results (i.e.UR comparison between LVIS and COCO prompt and qualitative results in Table.1(d)), we investigate that the unknown annotations in the test set are not adequate and more corresponding to the original 80 categories in COCO dataset. Therefore, existing “unknown recall” can not evaluate the detecting ability for unknown open-world objects beyond COCO annotation. To evaluate that, we use SAM(SAM) to generate annotations for all areas of similar objects in the test set. Though these annotations are noisy and contain many non-object box annotations, the recall for these can evaluate the unknown detection ability of detectors. The comparison between LVIS and COCO prompts shows that less distilled knowledge reduces the impact on learning known objects. In addition, less distilled knowledge impacts the performance of the knowledge distillation framework. As shown in OWOD split of COCO prompt, our detector does not learn better detecting performance for the original unknown objects. The results in LVIS w/o COCO prompt demonstrate our framework letter more unknown open-world objects. Through we exclude all corresponding coco categories from the LVIS prompt, our detector’s detection ability for original unknown objects does not seem to be impacted.
Ablating different detection transformer structures: As shown in Table.1(b), we compare our structure with original parallel structure (Zhu et al. 2020) and structure of (Ma et al. 2023). The results demonstrate that our structure has a better ability for alleviating the influence of distilled knowledge on known objects.
@@
Ablating different teachers: In Table.1(c), we compare different teachers, i.e.GLIP(Li et al. 2022) and SAM (SAM). For SAM, due to the absence of supervision confidence, we can not leverage the down-weight loss function. Therefore, the known detection ability is severely reduced. In addition, due to the over-detection ability of SAM, the detector can not distill a better unknown detection ability.
Ablating the Pseudo Labels: We also ablate the pseudo labels by controlling the presence or absence of the box score branch. However, owing to the absence of effective metrics, existing metrics cannot demonstrate the difference. In terms of Unknown Recall, they almost have the same performance on UR and UR(sam) when the box score branch is removed (i.e.While it is possible to add a lot of unknown annotations after adding SAM-generated annotations, there are too many non-object annotations. Also, they can not evaluate the performance of unknown object detection in absolute terms). Therefore, we ablate it according to the several visualization results, as shown in Fig.2. When the pseudo labels are removed, the model misdetect many unknown objects. The comparison demonstrates that the pseudo labels can help the detector better know about the world.
| Task IDs | Task 1 | Task 2 | Task 3 | Task 4 | ||||||||||
| Unknown | Known | Unknown | Known | Unknown | Known | Known | ||||||||
| Recall | mAP() | Recall | mAP() | Recall | mAP() | mAP() | ||||||||
| Metrics | Split | () | Current | () | Previously | Current | Both | () | Previously | Current | Both | Previously | Current | Both |
| UC-OWOD | OWOD | 2.4 | 50.7 | 3.4 | 33.1 | 30.5 | 31.8 | 8.7 | 28.8 | 16.3 | 24.6 | 25.6 | 15.9 | 23.2 |
| ORE-EBUI | 4.9 | 56.0 | 2.9 | 52.7 | 26.0 | 39.4 | 3.9 | 38.2 | 12.7 | 29.7 | 29.6 | 12.4 | 25.3 | |
| OW-DETR | 7.5 | 59.2 | 6.2 | 53.6 | 33.5 | 42.9 | 5.7 | 38.3 | 15.8 | 30.8 | 31.4 | 17.1 | 27.8 | |
| OCPL | 8.3 | 56.6 | 7.7 | 50.6 | 27.5 | 39.1 | 11.9 | 38.7 | 14.7 | 30.7 | 30.7 | 14.4 | 26.7 | |
| 2B-OCD | 12.1 | 56.4 | 9.4 | 51.6 | 25.3 | 38.5 | 11.6 | 37.2 | 13.2 | 29.2 | 30.0 | 13.3 | 25.8 | |
| Ours | 39.0 | 56.8 | 36.7 | 52.3 | 28.3 | 40.3 | 36.1 | 36.9 | 16.4 | 30.1 | 31.0 | 14.7 | 26.9 | |
| ORE-EBUI | MS-COCO | 1.5 | 61.4 | 3.9 | 56.5 | 26.1 | 40.6 | 3.6 | 38.7 | 23.7 | 33.7 | 33.6 | 26.3 | 31.8 |
| OW-DETR | 5.7 | 71.5 | 6.2 | 62.8 | 27.5 | 43.8 | 6.9 | 45.2 | 24.9 | 38.5 | 38.2 | 28.1 | 33.1 | |
| Ours | 60.9 | 69.4 | 60.0 | 63.8 | 26.9 | 44.4 | 58.6 | 46.2 | 28.0 | 40.1 | 41.8 | 29.6 | 38.7 | |
| Method | Settings | ||
| 10+10 | 15+5 | 19+1 | |
| ILOD | 63.2 | 65.8 | 68.2 |
| Faster ILOD | 62.1 | 67.9 | 68.5 |
| ORE | 64.5 | 68.5 | 68.8 |
| OW-DETR | 65.7 | 69.4 | 70.2 |
| Ours | 68.6 | 72.1 | 72.9 |
More Detailed Comparison With the Teacher
In order to keep the model’s ability to generalize to open-world objects in the comparison, we do not finetune the teacher on the known dataset. Because the finetuning process severely destroy the teacher’s generality. Although we also show the detection accuracy for the known classes, we only need to focus on the comparison of the detection ability for unknown objects. The results compared with the large pre-trained vision-language grounding model (GLIP) for OWOD problem are shown in Table.2. Regarding the number of parameters and FLOPs, our model is significantly smaller than GLIP. In particular, the inference speed of ours is of GLIP’s. Furthermore, the number of GLIP’s training data is 64M images, while our model only needs a small amount of data in each task of different splits almost of its. Compared with GLIP’s U-Recall of 37.0, 35.5, and 34.9 on Task 1, 2, and 3 of OWOD split, ours achieves 39.0, 36.7 and 36.1 in the corresponding tasks, achieving significant absolute gains up to 2.0, 1.2 and 1.2, respectively. In the MS-COCO split, our model achieves more than 5 point improvement in terms of U-Recall. This demonstrates that our model has a better ability to detect unknown objects for OWOD.
Qualitative Results: To present the ability to detect unknown objects intuitively, we select several open scenes from games and comics. These scenes contain the object out of the categories of LVIS (Gupta, Dollar, and Girshick 2019) text prompts. visualization results are shown in Fig.3. Benefits from the box score branch and pseudo labels, our detector evolves the novel unknown objects beyond the large pre-trained vision-language grounding model.
Comparison With SOTA OWOD Methods
@@
OWOD SPLIT: The results compared with the state-of-the-art methods on OWOD split for OWOD problem are shown in Table.3 (a). The ability of our model to detect unknown objects quantified by U-Recall is more than of those reported in previous state-of-the-art OWOD methods. Compared with the model 2B-OCD (Wu et al. 2022a) with the highest U-Recall of 12.1, 9.4 and 11.6 on Task 1, 2, and 3, ours achieves 39.0, 36.7, and 36.1 in the corresponding tasks, achieving significant absolute gains up to 26.9, 27.3 and 24.5, respectively. Benefits from the cascade decoder structure and down-weight training loss which mitigate the effect of unknown objects on detecting known objects, our model’s performance on known objects is also superior to most state-of-the-art methods.
@@
MS-COCO SPLIT: We report the results on the MS-COCO split in Table.3 (a). MS-COCO split mitigates data leakage across tasks and assigns more data to each Task, while our model receives a more significant boost compared with the OWOD split. Our model’s unknown object detection capability, quantified by U-Recall, is almost of those reported in previous state-of-the-art OWOD methods. Compared with OW-DETR’s U-Recall of 5.7, 6.2, and 6.9 on Task 1, 2, and 3, ours achieves 60.9, 60.0, and 58.6 in the corresponding tasks, achieving significant absolute gains up to 55.2, 53.8 and 51.7, respectively. This demonstrates that our model has the more powerful ability to retrieve new knowledge and detect unknown objects when faced with more difficult tasks.
@@
Incremental Object Detection: To intuitively present our detector’s ability for detecting object instances, we compare it to (Shmelkov, Schmid, and Alahari 2017; Peng, Zhao, and Lovell 2020; Joseph et al. 2021; Gupta et al. 2022) on the incremental object detection (IOD) task. We do not use assistance from the large pre-trained vision-language grounding model. We evaluate the experiments on three standard settings, where a group of classes (10, 5, and last class) is introduced incrementally to a detector trained on the remaining classes (10, 15, and 19), based on the PASCAL VOC 2007 dataset (Everingham et al. 2010). As the results shown in Table.3 (b), our model outperforms the existing method in a great migration on all three settings, indicating the power of cascade detection transformer for IOD.
Conclusions

### v2

@@ 6.4 Ablation Study
![[Uncaptioned image]](extracted/2312.08653v2/figure/rebuttual.png)
6.4.1 Ablating components
@@
To study the contribution of each component, we design ablation experiments in TABLE.IV. We set GLIP as our baseline, the Distillation improves the performance on unknown object detection but reduces the detector’s ability for known objects. Adding the down-weight training loss function significantly improves the performance on detecting known objects and incremental object detecting, achieving significant absolute gains up to more than 10 points. As we have analyzed, the cascade decoupling structure alleviates the inclusion of unknown objects on the known detecting performance and the confusion between the categories and locations of the same objects. It significantly improves the performance of detecting known objects with absolute gains up to more than 10 points, too. What’s more, these two combine effectively, improving performance with absolute gains up to 15.7, 21.1, and 16.0, without reducing the ability to detect unknown objects. Thus, each component has a critical role to play in open-world object detection.
Task IDs | Task 1 | Task 2 | Task 3 | Task 4 | |||||||||
Unknown | Known | Unknown | Known | Unknown | Known | Known | |||||||
Recall | mAP() | Recall | mAP() | Recall | mAP() | mAP() | |||||||
Metrics | () | Current | () | Previously | Current | Both | () | Previously | Current | Both | Previously | Current | Both |
UC-OWOD[37] | 2.4 | 50.7 | 3.4 | 33.1 | 30.5 | 31.8 | 8.7 | 28.8 | 16.3 | 24.6 | 25.6 | 15.9 | 23.2 |
ORE-EBUI[12] | 4.9 | 56.0 | 2.9 | 52.7 | 26.0 | 39.4 | 3.9 | 38.2 | 12.7 | 29.7 | 29.6 | 12.4 | 25.3 |
OW-DETR[13] | 7.5 | 59.2 | 6.2 | 53.6 | 33.5 | 42.9 | 5.7 | 38.3 | 15.8 | 30.8 | 31.4 | 17.1 | 27.8 |
OCPL[14] | 8.3 | 56.6 | 7.7 | 50.6 | 27.5 | 39.1 | 11.9 | 38.7 | 14.7 | 30.7 | 30.7 | 14.4 | 26.7 |
2B-OCD[15] | 12.1 | 56.4 | 9.4 | 51.6 | 25.3 | 38.5 | 11.6 | 37.2 | 13.2 | 29.2 | 30.0 | 13.3 | 25.8 |
Ours | 39.0 | 56.8 | 36.7 | 52.3 | 28.3 | 40.3 | 36.1 | 36.9 | 16.4 | 30.1 | 31.0 | 14.7 | 26.9 |
Task IDs | Task 1 | Task 2 | Task 3 | Task 4 | |||||||||
Unknown | Known | Unknown | Known | Unknown | Known | Known | |||||||
Recall | mAP() | Recall | mAP() | Recall | mAP() | mAP() | |||||||
Metrics | () | Current | () | Previously | Current | Both | () | Previously | Current | Both | Previously | Current | Both |
ORE-EBUI[12] | 1.5 | 61.4 | 3.9 | 56.5 | 26.1 | 40.6 | 3.6 | 38.7 | 23.7 | 33.7 | 33.6 | 26.3 | 31.8 |
OW-DETR[13] | 5.7 | 71.5 | 6.2 | 62.8 | 27.5 | 43.8 | 6.9 | 45.2 | 24.9 | 38.5 | 38.2 | 28.1 | 33.1 |
Ours | 60.9 | 69.4 | 60.0 | 63.8 | 26.9 | 44.4 | 58.6 | 46.2 | 28.0 | 40.1 | 41.8 | 29.6 | 38.7 |
6.4.2 Ablating different prompts
@@
To ablate the impact of text prompts for knowledge distillation, we do experiments on different text prompts, i.e.LVIS, LVIS without COCO, and COCO as shown in TABLE.V(a). From the experiment results (i.e.UR comparison between LVIS and COCO prompt and qualitative results in TABLE.V(d)), we investigate that the unknown annotations in the test set are not adequate and more corresponding to the original 80 categories in COCO dataset. Therefore, existing “unknown recall” can not evaluate the detecting ability for unknown open-world objects beyond COCO annotation. To evaluate that, we use SAM[58] to generate annotations for all areas of similar objects in the test set. Though these annotations are noisy and contain many non-object box annotations, the recall for these can evaluate the unknown detection ability of detectors. The comparison between LVIS and COCO prompts shows that less distilled knowledge reduces the impact on learning known objects. In addition, less distilled knowledge impacts the performance of the knowledge distillation framework. As shown in OWOD split of COCO prompt, our detector does not learn better detecting performance for the original unknown objects. The results in LVIS w/o COCO prompt demonstrate our framework letter more unknown open-world objects. Although we exclude all corresponding coco categories from the LVIS prompt, our detector’s detection ability for original unknown objects does not seem to be impacted.
6.4.3 Ablating different detection transformer structures
As shown in TABLE.V(b), we compare our structure with original parallel structure [2] and structure of [36]. CAT proposes the cascade decoupled decoding way that uses the shared decoder to decode twice for localization and identification, respectively. Unlike them, we directly employ two separate decoders to decouple the decoding process, and we specifically train these two decoders for localization and recognition, respectively. The results demonstrate that our structure has a better ability to alleviate the influence of distilled knowledge on known objects. Our decoupled architecture outperforms the decoupled decoding way of CAT[36] by 1.5 points in the known mAP. Moreover, our architecture surpasses the conventional DDETR [2] structure by 3.3 points.
6.4.4 Ablating different teachers
@@
In TABLE.V(c), we compare different teachers, i.e.GLIP[26] and SAM [58]. When distilling open-world knowledge from SAM, we leverage the “Segment Anything” that uses the predefined grid points and selecting strategy to generate pseudo labels. For SAM, due to the absence of supervision confidence, we can not leverage the down-weight loss function. Therefore, the known detection ability is severely reduced. In addition, due to the over-detection ability of SAM, the detector can not distill a better unknown detection ability.
6.5 Comparison With SOTA OWOD Methods
In this section, we conduct a detailed comparison with existing state-of-the-art methods on two existing benchmarks: OWOD and MS-COCO SPLIT, as well as on two benchmarks we propose: IntensiveSet and StandardSet. In addition, we compare our model with existing methods on the Incremental Object Detection task.
6.5.1 OWOD SPLIT
@@
The results compared with the state-of-the-art methods on OWOD split for OWOD problem are shown in TABLE.VI. The ability of our model to detect unknown objects quantified by U-Recall is more than of those reported in previous state-of-the-art OWOD methods. Compared with the model 2B-OCD [15] with the highest U-Recall of 12.1, 9.4 and 11.6 on Task 1, 2, and 3, ours achieves 39.0, 36.7, and 36.1 in the corresponding tasks, achieving significant absolute gains up to 26.9, 27.3 and 24.5, respectively. Benefits from the cascade decoder structure and down-weight training loss which mitigate the effect of unknown objects on detecting known objects, our model’s performance on known objects is also superior to most state-of-the-art methods.
6.5.2 MS-COCO SPLIT
@@
We report the results on the MS-COCO split in TABLE.VII. MS-COCO split mitigates data leakage across tasks and assigns more data to each Task, while our model receives a more significant boost compared with the OWOD split. Our model’s unknown object detection capability, quantified by U-Recall, is almost of those reported in previous state-of-the-art OWOD methods. Compared with OW-DETR’s U-Recall of 5.7, 6.2, and 6.9 on Task 1, 2, and 3, ours achieves 60.9, 60.0, and 58.6 in the corresponding tasks, achieving significant absolute gains up to 55.2, 53.8 and 51.7, respectively. This demonstrates that our model has the more powerful ability to retrieve new knowledge and detect unknown objects when faced with more difficult tasks.
Dataset | IntensiveSet | StandardSet | |||||||
Group | Metrics | Known | Unknown() | Known | Unknown() | ||||
mAP() | Recall | Precision | mAP | mAP() | Recall | Precision | mAP | ||
(I) | OWDETR[13] | 36.9 | 5.6 | 7.3 | 1.2 | 54.6 | 16.9 | 2.3 | 0.6 |
CAT[36] | 38.6 | 17.4 | 9.5 | 7.6 | 55.4 | 48.8 | 2.7 | 4.3 | |
UnSniffer[38] | 39.5 | 9.4 | 22.2 | 2.0 | 53.7 | 34.6 | 21.5 | 12.5 | |
UnSniffer[38] | 39.5 | 12.4 | 12.0 | 2.6 | 53.7 | 41.2 | 9.6 | 11.6 | |
(II) | SKDF-DW | 27.7 | 40.2 | 15.6 | 14.8 | 44.0 | 75.1 | 3.0 | 17.1 |
SKDF-CS | 30.7 | 39.7 | 16.2 | 16.8 | 48.3 | 75.2 | 3.1 | 9.2 | |
SKDF | 32.3 | 39.6 | 16.2 | 16.7 | 48.7 | 74.3 | 3.0 | 7.4 | |
SKDF | 32.3 | 35.8 | 37.9 | 24.5 | 48.7 | 66.8 | 11.4 | 24.4 |
6.5.3 Intensive Testing Set
The comparison results on IntensiveSet are shown in the left half of TABLE.VIII. The experimental results indicate that our model’s performance in detecting unknown objects surpasses that of existing state-of-the-art methods, achieving more than double the performance of the highest unknown recall for the category CAT [36], exceeding the highest unknown precision of Unsiffer [38] by 4.2 points, and surpassing the existing methods in unknown mAP by more than threefold. Meanwhile, our model is highly adaptable to unknown post-processing techniques. When incorporating the post-processing methods from UnSniffer [38], the performance of our model further increased by 21.7 points in unknown precision and 7.8 points in unknown mAP. When incorporating the post-processing, SKDF surpasses the UnSiffer in unknown precision by more than 15 points. However, although our component can mitigate the impact of unknown objects on the detection of known objects, there is still a certain gap between our model and the existing state-of-the-art models.
6.5.4 Standard Testing Set
In the right half of the TABLE.VIII, we compare SKDF with the sota methods on StandardSet. Our model surpasses the existing detection models in the vast majority of evaluation metrics, except for the detection performance of known objects and unknown precision. For the unknown precision, UnSiffer outperforms our SKDF. This is because the StandardSet primarily comprises common categories from COCO that are beyond the predefined classes of VOC, and UnSiffer is specifically tailored to work with these category distributions. However, this indeed reflects the shortcomings of our model. Particularly, when comparing the performance of our model on the IntensiveSet and StandardSet, it is evident that although our model has very good recall performance for unknown objects, its ability to recognize unknown objects still needs to be improved.
6.5.5 Incremental Object Detection
@@
To intuitively present our detector’s ability for detecting object instances, we compare it to [59, 60, 12, 13] on the incremental object detection (IOD) task. We do not use assistance from the large pre-trained vision-language grounding model. We evaluate the experiments on three standard settings, where a group of classes (10, 5, and last class) is introduced incrementally to a detector trained on the remaining classes (10, 15, and 19), based on the PASCAL VOC 2007 dataset [25]. As the results shown in TABLE.IX (b), our model outperforms the existing method in a great migration on all three settings, indicating the power of cascade detection transformer for IOD.
10 + 10 Setting | aero | cycle | bird | boat | bottle | bus | car | cat | chair | cow | table | dog | horse | bike | person | plant | sheep | sofa | train | tv | mAP |
ILOD | 69.9 | 70.4 | 69.4 | 54.3 | 48 | 68.7 | 78.9 | 68.4 | 45.5 | 58.1 | 59.7 | 72.7 | 73.5 | 73.2 | 66.3 | 29.5 | 63.4 | 61.6 | 69.3 | 62.2 | 63.2 |
Faster ILOD | 72.8 | 75.7 | 71.2 | 60.5 | 61.7 | 70.4 | 83.3 | 76.6 | 53.1 | 72.3 | 36.7 | 70.9 | 66.8 | 67.6 | 66.1 | 24.7 | 63.1 | 48.1 | 57.1 | 43.6 | 62.1 |
ORE - (CC + EBUI) | 53.3 | 69.2 | 62.4 | 51.8 | 52.9 | 73.6 | 83.7 | 71.7 | 42.8 | 66.8 | 46.8 | 59.9 | 65.5 | 66.1 | 68.6 | 29.8 | 55.1 | 51.6 | 65.3 | 51.5 | 59.4 |
ORE - EBUI | 63.5 | 70.9 | 58.9 | 42.9 | 34.1 | 76.2 | 80.7 | 76.3 | 34.1 | 66.1 | 56.1 | 70.4 | 80.2 | 72.3 | 81.8 | 42.7 | 71.6 | 68.1 | 77 | 67.7 | 64.5 |
OW - DETR | 75.4 | 63.9 | 57.9 | 50.0 | 52.0 | 70.9 | 79.5 | 72.4 | 44.3 | 57.9 | 59.7 | 73.5 | 77.7 | 75.2 | 76.2 | 44.9 | 68.8 | 65.4 | 79.3 | 69.0 | 65.7 |
Ours | 77.1 | 72.3 | 74.5 | 53.4 | 57.4 | 78.1 | 78.7 | 83.9 | 46.2 | 71.4 | 59.5 | 77.4 | 73.3 | 76.6 | 73.3 | 39.7 | 70.6 | 59.0 | 78.4 | 70.9 | 68.6 |
15 + 5 Setting | aero | cycle | bird | boat | bottle | bus | car | cat | chair | cow | table | dog | horse | bike | person | plant | sheep | sofa | train | tv | mAP |
ILOD | 70.5 | 79.2 | 68.8 | 59.1 | 53.2 | 75.4 | 79.4 | 78.8 | 46.6 | 59.4 | 59 | 75.8 | 71.8 | 78.6 | 69.6 | 33.7 | 61.5 | 63.1 | 71.7 | 62.2 | 65.8 |
Faster ILOD | 66.5 | 78.1 | 71.8 | 54.6 | 61.4 | 68.4 | 82.6 | 82.7 | 52.1 | 74.3 | 63.1 | 78.6 | 80.5 | 78.4 | 80.4 | 36.7 | 61.7 | 59.3 | 67.9 | 59.1 | 67.9 |
ORE - (CC + EBUI) | 65.1 | 74.6 | 57.9 | 39.5 | 36.7 | 75.1 | 80 | 73.3 | 37.1 | 69.8 | 48.8 | 69 | 77.5 | 72.8 | 76.5 | 34.4 | 62.6 | 56.5 | 80.3 | 65.7 | 62.6 |
ORE - EBUI | 75.4 | 81 | 67.1 | 51.9 | 55.7 | 77.2 | 85.6 | 81.7 | 46.1 | 76.2 | 55.4 | 76.7 | 86.2 | 78.5 | 82.1 | 32.8 | 63.6 | 54.7 | 77.7 | 64.6 | 68.5 |
OW - DETR | 78.0 | 80.7 | 79.4 | 70.4 | 58.8 | 65.1 | 84.0 | 86.2 | 56.5 | 76.7 | 62.4 | 84.8 | 85.0 | 81.8 | 81.0 | 34.3 | 48.2 | 57.9 | 62.0 | 57.0 | 69.4 |
Ours | 79.5 | 85.1 | 83.1 | 73.1 | 62.5 | 68.7 | 83.0 | 88.4 | 55.5 | 78.3 | 69.7 | 83.0 | 86.6 | 73.2 | 78.8 | 30.8 | 67.6 | 60.8 | 76.0 | 58.7 | 72.1 |
19 + 1 Setting | aero | cycle | bird | boat | bottle | bus | car | cat | chair | cow | table | dog | horse | bike | person | plant | sheep | sofa | train | tv | mAP |
ILOD | 69.4 | 79.3 | 69.5 | 57.4 | 45.4 | 78.4 | 79.1 | 80.5 | 45.7 | 76.3 | 64.8 | 77.2 | 80.8 | 77.5 | 70.1 | 42.3 | 67.5 | 64.4 | 76.7 | 62.7 | 68.2 |
Faster ILOD | 64.2 | 74.7 | 73.2 | 55.5 | 53.7 | 70.8 | 82.9 | 82.6 | 51.6 | 79.7 | 58.7 | 78.8 | 81.8 | 75.3 | 77.4 | 43.1 | 73.8 | 61.7 | 69.8 | 61.1 | 68.5 |
ORE - (CC + EBUI) | 60.7 | 78.6 | 61.8 | 45 | 43.2 | 75.1 | 82.5 | 75.5 | 42.4 | 75.1 | 56.7 | 72.9 | 80.8 | 75.4 | 77.7 | 37.8 | 72.3 | 64.5 | 70.7 | 49.9 | 64.9 |
ORE - EBUI | 67.3 | 76.8 | 60 | 48.4 | 58.8 | 81.1 | 86.5 | 75.8 | 41.5 | 79.6 | 54.6 | 72.8 | 85.9 | 81.7 | 82.4 | 44.8 | 75.8 | 68.2 | 75.7 | 60.1 | 68.8 |
OW - DETR | 82.2 | 80.7 | 73.9 | 56.0 | 58.6 | 72.1 | 82.4 | 79.6 | 48.0 | 72.8 | 64.2 | 83.3 | 83.1 | 82.3 | 78.6 | 42.1 | 65.5 | 55.4 | 82.9 | 60.1 | 70.2 |
Ours | 83.6 | 85.7 | 77.1 | 61.5 | 58.9 | 74.3 | 86.3 | 81.5 | 52.2 | 78.4 | 71.4 | 81.9 | 84.6 | 80.2 | 80.8 | 39.9 | 68.3 | 63.3 | 84.6 | 63.0 | 72.9 |
6.5.6 Qualitative Comparison with OWOD method
We exhibit more qualitative results in Fig.6 and compare them to the state-of-the-art OWOD detectors. The results show that our model can detect most of the potentially unknown open-world objects in the scene, far exceeding the effectiveness of existing models. We define the 20 categories of the PASCAL VOC [25] dataset as known classes, and all categories outside of these are considered unknown classes. As shown in the first column, our model accurately detects all the foreground categories in the scene and identifies everything outside of the predefined known classes as unknown, such as the books on the table and the teddy bear, which were not detected by the state-of-the-art model. As shown in the second column, our model can accurately detect even objects such as the picture frames on the wall. In the third column, the state-of-the-art model mistakenly identifies the characters in the background as unknown objects, but our model is not confused by this. In the comparison of the last column, our model also greatly surpasses the detection performance of the state-of-the-art model.
7 Societal Impact, Limitations and Future Works
Open-world object detection makes artificial intelligence smarter to face more problems in real life. It takes object detection to a cognitive level, as the model requires more than simply remembering the objects learned, it requires deeper thinking about the scene. When it comes to applications like autonomous driving, the significance of open-world object detection becomes even more pronounced. In such scenarios, vehicles need to rapidly and accurately identify and comprehend various objects and obstacles on the road, including but not limited to pedestrians, other vehicles, traffic signals, and road signs. The breakthrough in open-world object detection will render autonomous driving systems more intelligent as they can handle unforeseen or rare situations, not limited to pre-trained object categories.
Although our results demonstrate significant improvements over existing state-of-the-art methods, the performances are still on the lower side due to the challenging nature of the open-world detection problem. In this paper, we are mainly committed to enhancing the model’s ability to explore unknown classes. However, the confidence level and recognition ability of our model for the detection of unknown objects still need to be improved, and this is what we will strive for in the future. Currently, our model still faces the issue of mistakenly detecting the background as unknown objects, and the benchmark we proposed is not yet comprehensive, including only a single-moment task.
In our future work, these two points will be the main focus of our research. In addition, post-processing operations for prediction boxes of unknown objects are urgently needed, so developing a post-processing operation similar to NMS, dedicated to unknown objects, is also a direction that requires our research. To facilitate the integration of open-world object detection algorithms into everyday use and contribute positively to society, we are fully committed to diligently pursuing this goal with great care and effort.
8 Conclusions